# Subsurface Scattering for 3D Gaussian Splatting

### v1

@@ 2 Related Work
There have been a multitude of prior works in SSS. Jensen et al. [16] presents a practical model for subsurface light transport in translucent materials that captures effects beyond BRDF models. Donner et al. [9] further explore light diffusion in multi-layered translucent materials using a variant of the Kubelka-Munk theory. Lensch et al. [22] introduce a rendering method for translucent diffuse objects, in which viewpoint and illumination can be modified at interactive rates. and filtered using the precomputed kernels. Vicini et al. [36] introduce a new shape-adaptive BSSRDF model that is based on a conditional variational autoencoder which learns to sample from a reference distribution produced by a brute-force volumetric path tracer. The distribution is conditional on both material properties and a set of features characterizing geometric variations in the neighborhood of the incident location. Zheng et al.[47] learn neural representations for participating media with a complete simulation of global illumination. They estimate direct illumination via ray tracing and compute indirect illumination with Spherical Harmonics. Zhu et al. [48] propose a novel framework for learning the radiance transfer field via volume rendering and utilizing various appearance cues to refine the geometry end-to-end. They extend relighting and reconstruction prospects to tackle a wider range of materials in a data-driven manner. They use a NeRF-like architecture to represent subsurface scattering based on known incident light directions from controlled acquisition. Similarly, Neural Radiance Transfer Fields [25] use one light at a time (OLAT) data as supervision to learn global light transport. This data representation is very similar to ours, however, their dataset of translucent objects has not been released so far. Due to using a NeRF-only approach they lack editing capabilities and have longer run times. Object-Centric Neural Scattering Functions (OSF) [45] introduce a framework for representing and rendering objects under varying lighting conditions and from arbitrary viewpoints using object-centric neural scattering functions. OSFs allow for flexible composition of scenes with multiple objects, each maintaining realistic interactions with light and shadows but they also suffer from long training time, even their variant based on KiloNeRF [31].
@@ 3.2 Subsurface Scattering for 3D Gaussian Splatting
where $f_{\Theta}$ is the neural network, $\mu$ and $\Sigma$ are the mean and covariance of a 3D Gaussian, $n$ is the normal, $\omega_{\text{in}}$ and $\omega_{\text{out}}$ are the incident and outgoing light directions, and $v$ is the shadowing term (see below). We apply a Fourier encoding to $\mu$ as in [27] which is omitted in Equation 6. The network is not only estimating the outgoing SSS radiance $L_{\text{out (SSS)}}$ but at the same time the incident light $\bar{L}_{\text{in}}$ is predicted which will later be used to evaluate the direct reflection at $\mu$. We want to note that $\bar{L}_{\text{in}}$ is a prediction of our model that is close to the true physical quantity $L_{\text{in}}$ but might be slightly offset to compensate for limitations in the model. By forcing the network to predict the incident as well as the outgoing SSS radiance we implicitly force the network to learn about the local and global light transport in the scene. $L_{\text{out (SSS)}}$ and $\bar{L}_{\text{in}}$ predictions share the same MLP but apply separate output heads (Fig. 2).
@@
With this formulation, we achieve joint learning of geometry and the direct and global appearance of SSS objects. See Appendix G for more details on the architecture and the training. Further, the choice of the 3D GS representation and our lightweight MLP combined with explicit PBR shading allow for real-time rendering speeds. Due to the deferred shading approach, we can capture high-frequency details such as specular highlights. Furthermore, the decomposition into explicit distinct appearance components allows for a high degree of editability, even controlling the degree of SSS effects.
@@ 4.1 Experimental setup
SSS GS builds upon the 3D GS framework [18] and its extension Relightable 3D Gaussians (R3DGS) [10] that we in turn extend to capture the subsurface scattering effects of an object. For our implicit representation of the scattering component and the incident light prediction, we use a shallow MLP with $3$ layers with Leaky-ReLU activations [11]. The whole pipeline is implemented with Pytorch using the provided custom CUDA kernels from [19, 10] for rendering. For more details on the training setup see Appendix G.
@@ 4.2 Qualitative Results
Our qualitative results are shown in Figure 3 and Figure 5. However, relighting and novel view synthesis are best experienced in the supplementary video. In 3 we showcase results rendered from our test set of objects from the synthetic (first two rows) and real-world (bottom two rows) part of our newly created dataset. In addition to novel view synthesis (view 0, view 1) we can freely change the light direction which we show for both views. We present the decomposition of the object into the surface appearance parameters, SSS effects and incident light map as they are used to render the final views on the left. Our method captures the SSS effect well using the SSS residual and adds the object’s specularity from the PBR shading. Together with the predicted incident light the model achieves a plausible decomposition into basecolor, roughness and metalness material parameters. Note how the volumetric component can represent the complex light transport inside the entire volume. This is particularly observable in the case of the Tupperware object while the normals show a detailed representation of the surface. Find further quantitative analysis regarding the decomposition in section F.
@@
In Figure 4 we show that Relightable 3D Gaussian (R3DGS) [10] fail to capture the subsurface scattering effects of the object. As they only allow training on a static scene without any dynamic lighting the scattering component is baked into the basecolor which will fail to represent the SSS for a new light position (see also Sec. C) Our method can capture the subsurface scattering effects and the specular highlights of the object. Acheiving clearer results than R3DGS [10] in that regard due to the formulation of shading in image space. Figure 4 also visualizes the difference between shading in world space vs. our deferred shading approach in image space. As visualized on the left the Gaussians occupy different, sometimes very large areas along the surface limiting the rendering of high-frequency effects like specular highlights. NeRF-based methods have a lot more training capacity due to their large MLPs that can represent the complex light transport connected to SSS. For a fair comparison we select KiloOSF a variant of OSF [45] that is optimized for real-time rendering. Compared to KiloOSF [45] in Figure 5 it becomes apparent that also a small MLP paired with our explicit shading approach can achieve higher quality results after a shorter training time. Even after 20 hours of training, there are still some artifacts visible in the geometry that stem from the voxelization performed by the underlying KiloNeRF [31]. While the Stanford Bunny object overall works well, the soap bar and car toy object are lacking in the representation of the surface reflectivity and show frequent errors in the geometry.
@@ 4.4 Applications
Find details for further qualitative and quantitative analysis of the application possibilities of our method such as editing applications (Sec. E) and image based relighting possibilities (Sec. D) in the appendix.
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP 02, project number: 276693517. This work was supported by the Tübingen AI Center. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan-Niklas Dihlmann and Arjun Majumdar. In cooperation with Sony Germany, the authors would like to thank Mr. Zoltan Facius and Mr. Prasan Ashok Shedligeri.

### v2

@@ 2 Related Work
There have been a multitude of prior works in SSS. Jensen et al. [16] presents a practical model for subsurface light transport in translucent materials that captures effects beyond BRDF models. Donner et al. [9] further explore light diffusion in multi-layered translucent materials using a variant of the Kubelka-Munk theory. Lensch et al. [22] introduce a rendering method for translucent diffuse objects, in which viewpoint and illumination can be modified at interactive rates. and filtered using the precomputed kernels. Vicini et al. [37] introduce a new shape-adaptive BSSRDF model that is based on a conditional variational autoencoder which learns to sample from a reference distribution produced by a brute-force volumetric path tracer. The distribution is conditional on both material properties and a set of features characterizing geometric variations in the neighborhood of the incident location. Zheng et al.[48] learn neural representations for participating media with a complete simulation of global illumination. They estimate direct illumination via ray tracing and compute indirect illumination with Spherical Harmonics. Zhu et al. [49] propose a novel framework for learning the radiance transfer field via volume rendering and utilizing various appearance cues to refine the geometry end-to-end. They extend relighting and reconstruction prospects to tackle a wider range of materials in a data-driven manner. They use a NeRF-like architecture to represent subsurface scattering based on known incident light directions from controlled acquisition. Similarly, Neural Radiance Transfer Fields [25] use one light at a time (OLAT) data as supervision to learn global light transport. This data representation is very similar to ours, however, their dataset of translucent objects has not been released so far. Due to using a NeRF-only approach they lack editing capabilities and have longer run times. Object-Centric Neural Scattering Functions (OSF) [46] introduce a framework for representing and rendering objects under varying lighting conditions and from arbitrary viewpoints using object-centric neural scattering functions. OSFs allow for flexible composition of scenes with multiple objects, each maintaining realistic interactions with light and shadows but they also suffer from long training time, even their variant based on KiloNeRF [31]. Further, in the realm of Gaussian Splatting, there have been works on human avatars incorperating some modeling of subsurface scattering with Spherical Harmonics [32].
@@ 3.2 Subsurface Scattering for 3D Gaussian Splatting
where $f_{\Theta}$ is the neural network, $\mu$ and $\Sigma$ are the mean and covariance of a 3D Gaussian, $n$ is the normal, $\omega_{\text{in}}$ and $\omega_{\text{out}}$ are the incident and outgoing light directions, and $v$ is the shadowing term (see below). We apply a Fourier encoding to $\mu$ as in [27] which is omitted in Equation 6. The network is not only estimating the outgoing SSS radiance $L_{\text{out (SSS)}}$ but at the same time the incident light $\bar{L}_{\text{in}}$ is predicted which will later be used to evaluate the direct reflection at $\mu$. We want to note that $\bar{L}_{\text{in}}$ is a prediction of our model that is close to the true physical quantity $L_{\text{in}}$ but might be slightly offset to compensate for limitations in the model. By forcing the network to predict the incident as well as the outgoing SSS radiance we let the network implicitly learn about the local and global light transport in the scene. $L_{\text{out (SSS)}}$ and $\bar{L}_{\text{in}}$ predictions share the same MLP but apply separate output heads (Fig. 2).
@@
With this formulation, we achieve joint learning of geometry and the direct and global appearance of SSS objects. See Appendix F for more details on the architecture and the training. Further, the choice of the 3D GS representation and our lightweight MLP combined with explicit PBR shading allow for real-time rendering speeds. Due to the deferred shading approach, we can capture high-frequency details such as specular highlights. Furthermore, the decomposition into explicit distinct appearance components allows for a high degree of editability, even controlling the degree of SSS effects.
@@ 4.1 Experimental setup
SSS GS builds upon the 3D GS framework [18] and its extension Relightable 3D Gaussians (R3DGS) [10] that we in turn extend to capture the subsurface scattering effects of an object. For our implicit representation of the scattering component and the incident light prediction, we use a shallow MLP with $3$ layers with Leaky-ReLU activations [11]. The whole pipeline is implemented with Pytorch using the provided custom CUDA kernels from [19, 10] for rendering. For more details on the training setup see Appendix F.
@@ 4.2 Qualitative Results
Our qualitative results are shown in Figure 3 and Figure 6. However, relighting and novel view synthesis are best experienced in the supplementary video. In 3 we show results rendered from our test set of objects from the synthetic (first two rows) and real-world (bottom two rows) part of our newly created dataset. In addition to novel view synthesis (view 0, view 1) we can freely change the light direction which we show for both views. We present the decomposition of the object into the surface appearance parameters, SSS effects and incident light map as they are used to render the final views on the left. Our method captures the SSS effect well using the SSS residual and adds the object’s specularity from the PBR shading. Together with the predicted incident light the model achieves a plausible decomposition into basecolor, roughness and metalness material parameters. Note how the volumetric component can represent the complex light transport inside the entire volume. This is particularly observable in the case of the Tupperware object while the normals show a detailed representation of the surface. Find further quantitative analysis regarding the decomposition in section E.
@@
In Figure 5 we show that Relightable 3D Gaussian (R3DGS) [10] fail to capture the subsurface scattering effects of the object. As they only allow training on a static scene without any dynamic lighting the scattering component is baked into the basecolor which will fail to represent the SSS for a new light position (see also Sec. C) Our method can capture the subsurface scattering effects and the specular highlights of the object. Achieving clearer results than R3DGS [10] in that regard due to the formulation of shading in image space. Figure 5 also visualizes the difference between shading in world space vs. our deferred shading approach in image space. As visualized on the left the Gaussians occupy different, sometimes very large areas along the surface limiting the rendering of high-frequency effects like specular highlights. NeRF-based methods have a lot more training capacity due to their large MLPs that can represent the complex light transport connected to SSS. For a fair comparison we select KiloOSF a variant of OSF [46] that is optimized for real-time rendering. Compared to KiloOSF [46] in Figure 6 it becomes apparent that also a small MLP paired with our explicit shading approach can achieve higher quality results after a shorter training time. Even after 20 hours of training, there are still some artifacts visible in the geometry that stem from the voxelization performed by the underlying KiloNeRF [31]. While the Stanford Bunny object overall works well, the soap bar and car toy object are lacking in the representation of the surface reflectivity and show frequent errors in the geometry.
@@ 4.4 Applications
Find details on further qualitative and quantitative analysis of the application possibilities of our method such as image based relighting (Sec. D) in the appendix.
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP 02, project number: 276693517. This work was supported by the Tübingen AI Center. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan-Niklas Dihlmann and Arjun Majumdar. In cooperation with Sony Semiconductor Solutions -Europe, the authors would like to thank Mr. Zoltan Facius and Mr. Prasan Ashok Shedligeri.